# Revisiting Multimodal KV Cache Compression: A Frequency-Domain-Guided Outlier-KV-Aware Approach

### v1

@@ 5 Conclusion
[…]
@@
  • B: More Observations
    • B.1: Frequency-domain energy distribution of the KV matrices
    • B.2: Frequency-domain energy distribution of KV matrices across different layers
  • C: More Efficiency Analysis
  • D: Case Study
  • E: Benchmarks
    • E.1 Multi-Images Understanding Benchmarks
    • E.2 High Resolution Benchmarks
    • E.3 Video Understanding Benchmark
  • F: Models
    • F.1 Qwen2.5-VL-7B-Instruct
    • F.2 LLaVA-OneVision-1.5-8B-Instruct
  • G: Limitation and Future Work
@@ Appendix A More Experiments in different KV Retention Ratios
Appendix B More Observations
@@
B.1 Frequency-domain energy distribution of the KV matrices
We observe that the frequency-domain energy of KV matrices is predominantly concentrated at low frequency, with high frequency components occupying a relatively small proportion. To further validate this phenomenon, we conduct additional experiments across two models—Qwen2.5-VL-7B-Instruct [wang2024qwen2] and LLaVA-OneVision-1.5-8B-Instruct [LLaVAOneVision15] using eight datasets. Fig. 7 presents the results for Qwen2.5-VL-7B-Instruct, while Fig. 8 shows the results for LLaVA-OneVision-1.5-8B-Instruct. Combining the findings from both figures, we observe that the phenomenon of KV matrices frequency-domain energy concentration in the low-frequency range is significantly established.
B.2 Frequency-domain energy distribution of KV matrices across different layers
We observe clear differences in the frequency-domain energy distribution of KV matrices across layers. To further validate this phenomenon, we conduct additional experiments on two models—Qwen2.5-VL-7B-Instruct [wang2024qwen2] and LLaVA-OneVision-1.5-8B-Instruct [LLaVAOneVision15]—using eight datasets. The results for Qwen2.5-VL-7B-Instruct are shown in Fig. 9, and those for LLaVA-OneVision-1.5-8B-Instruct are presented in Fig. 10. Taken together, these results provide strong evidence that the frequency-domain energy distribution of KV matrices varies substantially across different layers.
Appendix C More Efficiency Analysis
@@
Appendix D Case Study
@@
To further demonstrate the advantages of our approach, we have selected several reasoning examples, as shown in Fig. 11. We compare FlashCache against state-of-the-art KV cache eviction methods: StreamingLLM [xiao2023efficient], H2O [zhang2023h2o], SnapKV [li2024snapkv], LOOK-M [wan2024look], and MEDA [wan2025meda]. We conduct comparisons across different image lengths, different text lengths, and different tasks. KV Cache retention ratio is setting as 0.05. Experimental results indicate that under high compression ratios, FlashCache preserves more of the model’s inherent performance characteristics. At the same time, we find that FlashCache enables the model to answer questions it previously could not. We believe FlashCache filters out noise in some input data, thereby improving the model’s performance under certain conditions.
Appendix E Benchmarks
E.1 Multi-Images Understanding Benchmarks
@@
MileBench[song2024milebench]. MileBench is a multimodal long-context benchmark designed to evaluate how well Multimodal Large Language Models handle complex multi-image inputs over extended contexts. It is built from 6,440 long-context samples collected from 29 publicly available or self-constructed datasets, with each sample containing multiple images and accompanying text. In our paper, we use four task types from MileBench: Task T (Temporal Multi-Image), which focuses on understanding and reasoning over sequences of images with temporal relationships; Task S (Semantic Multi-Image), which requires integrating complementary semantic information distributed across multiple images; NH (Needle in a Haystack), which tests the model’s ability to retrieve a small but crucial visual or textual detail from long, cluttered multimodal contexts; and IR (Image Retrieval), which assesses whether the model can accurately locate or identify the correct image given a textual description or multimodal query. Together, these tasks make MileBench a comprehensive tool for probing long-context multi-image comprehension, reasoning, and retrieval abilities.
@@
MUIRBench[wang2024muirbench]. MUIRBench is a comprehensive benchmark for robust multi-image understanding in multimodal large language models. It contains 11,264 images and 2,600 multiple-choice questions, covering 12 diverse multi-image understanding tasks (such as scene understanding, temporal ordering, and visual retrieval) and 10 categories of inter-image relations (including multiview, narrative, and complementary relations). Each example is constructed in a pairwise manner, where an answerable instance is paired with a minimally modified unanswerable variant, enabling robust evaluation of whether models truly integrate information across multiple images rather than relying on superficial cues. This design makes MUIRBench a strong stress test for multi-image reasoning and a useful complement to standard single-image vision–language benchmarks.
@@
MMMU[yue2023mmmu]. MMMU (Massive Multi-discipline Multimodal Understanding) is a large-scale benchmark designed to evaluate foundation models on expert-level, university and beyond problems across a wide range of real-world disciplines. It contains over 10,000 multimodal multiple-choice questions spanning fields such as medicine, law, physics, chemistry, engineering, and the humanities, with inputs that combine text, diagrams, charts, tables, and other images. Many questions involve not just a single image but multiple related figures or panels (for example several diagrams, tables, or subplots for one problem) that must be interpreted jointly, allowing the benchmark to test genuine multi-image reasoning rather than isolated image understanding. All items are carefully curated from real exam or professional materials, emphasizing high-level reasoning, domain knowledge, and cross-modal comprehension, and are widely used to assess how close multimodal large language models are to human expert performance in complex, knowledge-intensive scenarios.
E.2 High Resolution Benchmarks
@@
V*[vstar]. V* is an LLM-guided visual search mechanism designed to help multimodal large language models actively locate task-relevant regions in an image instead of passively encoding the whole frame. When paired with an MLLM, it underpins the SEAL (“Show, Search, and Tell”) meta-architecture, which iteratively queries the image to refine visual grounding and reasoning. To evaluate this capability, the authors introduce V* Bench, a dedicated benchmark built from 191 high-resolution natural images (average resolution around 2246×1582) that emphasize fine-grained details and visually crowded scenes. V* Bench focuses on detailed visual grounding through attribute recognition and spatial relationship reasoning questions, making it particularly suitable for testing whether models can reliably handle high-resolution inputs and answer questions about small or easily overlooked visual elements.
@@
HR-Bench[hrbench]. HR-Bench is a high-resolution visual question answering benchmark designed to assess the fine-grained perception ability of multimodal large language models on ultra high-resolution images. It consists of two sub-tasks: Fine-grained Single-instance Perception (FSP), which contains 100 samples focusing on detailed attribute recognition, OCR, and visually grounded prompting within a single image, and Fine-grained Cross-instance Perception (FCP), which contains 100 samples targeting map understanding, chart analysis, and spatial relationship reasoning across multiple visual elements. HR-Bench is provided in both 8K and 4K settings: the 8K version uses images with approximately 8K-level resolution, while the 4K version is obtained by carefully annotating objects in the 8K images and cropping them into 4K regions, enabling rigorous evaluation of whether models can maintain and exploit fine details under realistic high-resolution conditions.
E.3 Video Understanding Benchmark
@@
FAVOR-Bench[tu2025favor]. FAVOR-Bench is a comprehensive benchmark for fine-grained video motion understanding that targets the ability of multimodal large language models to perceive and reason about detailed temporal dynamics rather than just coarse scene changes. It contains 1,776 videos from both ego-centric and third-person perspectives, each accompanied by structured manual motion annotations. On top of this data, FAVOR-Bench provides both close-ended and open-ended evaluation settings: for close-ended evaluation, it offers 8,184 multiple-choice question–answer pairs organized into six motion-centric sub-tasks, while for open-ended evaluation it includes a GPT-assisted caption assessment protocol together with a cost-efficient LLM-free evaluation method. Through these tasks, FAVOR-Bench enables systematic measurement of how well models recognize, localize, and describe fine-grained motions in realistic videos, and serves as a challenging testbed for improving motion-sensitive video understanding.
Appendix F Models
F.1 Qwen2.5-VL-7B-Instruct
@@ Appendix F Models
Qwen2.5-VL-7B-Instruct [wang2024qwen2] is a 7-billion-parameter vision–language model from the Qwen2.5-VL family, instruction-tuned to handle a wide range of multimodal tasks with text, images, documents, charts, and video as input. It combines a strong language backbone with a visual encoder that supports dynamic resolutions and long-context processing, enabling capabilities such as document analysis, OCR, chart and layout understanding, temporal reasoning over videos, and structured output generation (for example JSON and bounding boxes for visual localization).
F.2 LLaVA-OneVision-1.5-8B-Instruct
@@
LLaVA-OneVision-1.5-8B-Instruct [LLaVAOneVision15] is an 8-billion-parameter large multimodal model in the LLaVA-OneVision-1.5 family, designed as a fully open, efficient framework for high-quality multimodal training. It is trained on native-resolution images and curated instruction data, aiming to deliver strong performance on diverse vision–language benchmarks while keeping computational costs relatively low. The model follows the LLaVA-style visual instruction tuning paradigm and is positioned as a general-purpose multimodal assistant capable of fine-grained image understanding, reasoning, and dialogue under an open-source training and evaluation pipeline.
Appendix G Limitation and Future Work
@@
Method LLaVA-OneVision-1.5-8B Qwen2.5-VL-7B Qwen2.5-VL-32B
Task T Task S NH IR Task T Task S NH IR Task T Task S NH IR
$\rho=0.1$
StreamingLLM [xiao2023efficient] 46.81 58.40 6.72 12.5 54.71 59.23 6.56 15.83 57 56.15 2.5 20.00
H2O [zhang2023h2o] 46.84 58.51 15.94 11.83 54.58 58.65 6.25 14.83 56.94 58.33 2.5 21.0
SnapKV [li2024snapkv] 46.97 57.62 10.47 12.67 55.58 61.30 7.19 15.50 57.06 58.71 16.06 18.83
LOOK-M [wan2024look] 46.84 58.51 15.94 11.83 55.08 58.54 5.78 13.67 51.91 58.99 2.81 24.67
MEDA [wan2025meda] 44.23 52.76 14.06 12.17 54.24 57.05 5.47 13.67 51.31 48.45 2.97 28.67
$\rho=0.05$
StreamingLLM [xiao2023efficient] 46.52 57.05 5.16 12.00 49.46 53.64 5.63 15.00 51.94 42.34 1.25 26.83
H2O [zhang2023h2o] 47.01 56.99 12.50 11.50 50.47 52.77 5.31 11.33 52.31 48.44 1.25 23.67
SnapKV [li2024snapkv] 45.97 57.82 7.97 13.00 51.52 56.33 6.25 15.33 55.69 49.07 6.25 19.83
LOOK-M [wan2024look] 48.01 56.99 12.50 11.50 51.69 53.67 5.31 16.33 49.97 45.77 2.5 27.83
MEDA [wan2025meda] 38.40 48.25 10.31 13.00 51.74 52.95 5.16 15.33 51.86 50.01 2.65 26.33

### v2

@@ 5 Conclusion
6 Acknowledgments
This work is supported by National Key Research and Development Program of China (No. 2022ZD0160101), Shanghai Natural Science Foundation (No. 23ZR1402900), Shanghai Science and Technology Commission Explorer Program Project (24TS1401300), Shanghai Municipal Science and Technology Major Project (No.2021SHZDZX0103). The computations in this research were performed using the CFFF platform of Fudan University.
@@ 6 Acknowledgments
  • B: More Experiments on Decoding Stage KV Retention
  • C: More Ablation Experiments on Low-Pass Cutoff Factor $\gamma$
  • D: More Observations
    • D.1: Frequency-domain energy distribution of the KV matrices
    • D.2: Frequency-domain energy distribution of KV matrices across different layers
    • D.3: The impact on model performance of prioritizing the removal of KV pairs under different removal ratios
  • E: More Efficiency Analysis
  • F: Case Study
  • G: Benchmarks
    • G.1 Multi-Images Understanding Benchmarks
    • G.2 High Resolution Benchmarks
    • G.3 Video Understanding Benchmark
  • H: Models
    • H.1 Qwen2.5-VL-7B-Instruct
    • H.2 LLaVA-OneVision-1.5-8B-Instruct
  • I: Limitation and Future Work
@@ Appendix A More Experiments in different KV Retention Ratios
Appendix B More Experiments on Decoding Stage KV Retention
Following established baselines (LOOK-M, MEDA), we focus on compressing prefill KV Cache while fully retaining decoding KVs. To validate robustness in long-generation scenarios, we evaluate FlashCache on the MathVista(COT) benchmark. As shown in Tab.9, FlashCache also maintains performance advantages in reasoning-heavy task.
Appendix C More Ablation Experiments on Low-Pass Cutoff Factor $\gamma$
Additional ablation on LLaVA-OneVision-8B (Tab.10) aligns with our Qwen results, identifying $\gamma\in[0.1,0.2]$ as optimal and robust. Theoretically, a lower $\gamma$ creates smoother Base KV that forces the deviation term to capture critical high-frequency features. Conversely, a high $\gamma$ causes Base KV to overfit, making outliers indistinguishable. Thus, $\gamma\in[0.1,0.2]$ is a robust guideline across models and datasets.
Method LLaVA-OneVision-1.5-8B Qwen2.5-VL-7B Qwen2.5-VL-32B
Task T Task S NH IR Task T Task S NH IR Task T Task S NH IR
$\rho=0.1$
StreamingLLM [41] 46.81 58.40 6.72 12.5 54.71 59.23 6.56 15.83 57 56.15 2.5 20.00
H2O [49] 46.84 58.51 15.94 11.83 54.58 58.65 6.25 14.83 56.94 58.33 2.5 21.0
SnapKV [21] 46.97 57.62 10.47 12.67 55.58 61.30 7.19 15.50 57.06 58.71 16.06 18.83
LOOK-M [35] 46.84 58.51 15.94 11.83 55.08 58.54 5.78 13.67 51.91 58.99 2.81 24.67
MEDA [34] 44.23 52.76 14.06 12.17 54.24 57.05 5.47 13.67 51.31 48.45 2.97 28.67
$\rho=0.05$
StreamingLLM [41] 46.52 57.05 5.16 12.00 49.46 53.64 5.63 15.00 51.94 42.34 1.25 26.83
H2O [49] 47.01 56.99 12.50 11.50 50.47 52.77 5.31 11.33 52.31 48.44 1.25 23.67
SnapKV [21] 45.97 57.82 7.97 13.00 51.52 56.33 6.25 15.33 55.69 49.07 6.25 19.83
LOOK-M [35] 48.01 56.99 12.50 11.50 51.69 53.67 5.31 16.33 49.97 45.77 2.5 27.83
MEDA [34] 38.40 48.25 10.31 13.00 51.74 52.95 5.16 15.33 51.86 50.01 2.65 26.33
Table 9: Comparison on MathVista (Qwen2.5-VL-7B, $\rho=0.1$).
Method Full Cache StreamLLM H2O SnapKV LOOK-M
MathVista 68.2 40.8 50.9 50.9 47.3
Method MEDA VLCache Quest PyramidKV FlashCache
MathVista 51.6 22.9 22.9 21.1 52.6
Table 10: Ablation Study on low-pass filter cutoff factor $\gamma$ on LLaVA-OneVision-8B.
$\gamma$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
INIAH 27.50 24.69 18.44 14.06 11.88 11.25 10.00 10.31 10.00
GPR1200 12.33 12.83 11.50 12.17 11.67 11.83 12.33 12.33 12.17
Appendix D More Observations
@@
D.1 Frequency-domain energy distribution of the KV matrices
We observe that the frequency-domain energy of KV matrices is predominantly concentrated at low frequency, with high frequency components occupying a relatively small proportion. To further validate this phenomenon, we conduct additional experiments across two models—Qwen2.5-VL-7B-Instruct [37] and LLaVA-OneVision-1.5-8B-Instruct [4] using eight datasets. Fig. 7 presents the results for Qwen2.5-VL-7B-Instruct, while Fig. 8 shows the results for LLaVA-OneVision-1.5-8B-Instruct. Combining the findings from both figures, we observe that the phenomenon of KV matrices frequency-domain energy concentration in the low-frequency range is significantly established.
D.2 Frequency-domain energy distribution of KV matrices across different layers
We observe clear differences in the frequency-domain energy distribution of KV matrices across layers. To further validate this phenomenon, we conduct additional experiments on two models—Qwen2.5-VL-7B-Instruct [37] and LLaVA-OneVision-1.5-8B-Instruct [4]—using eight datasets. The results for Qwen2.5-VL-7B-Instruct are shown in Fig. 9, and those for LLaVA-OneVision-1.5-8B-Instruct are presented in Fig. 10. Taken together, these results provide strong evidence that the frequency-domain energy distribution of KV matrices varies substantially across different layers.
Table 11: Method Latency in per layer with 32K/64K sequence length (metric: ms).
Method Prefilling H2O SnapKV LOOK-M MEDA FlashCache
32K 53.63 1297.27 76.62 1290.64 OOM 12.45
64K 148.08 OOM OOM OOM OOM 28.12
@@ D.2 Frequency-domain energy distribution of KV matrices across different layers
D.3 The impact on model performance of prioritizing the removal of KV pairs under different removal ratios
In Fig.11, the frequency-domain outlier phenomenon is consistent across diverse model architectures, parameter scales, and datasets. This evidence substantiates that our core insight is inherently general rather than model/dataset-specific.
Appendix E More Efficiency Analysis
@@
As shown in Tab.11, FlashCache’s overhead is negligible (12.45ms at 32K). Unlike baselines that suffer from OOM due to attention re-computation, FlashCache eliminates attention dependency, ensuring memory stability at 64K while introducing substantially lower prefilling overhead than existing methods.
Appendix F Case Study
@@
To further demonstrate the advantages of our approach, we have selected several reasoning examples, as shown in Fig. 12. We compare FlashCache against state-of-the-art KV cache eviction methods: StreamingLLM [41], H2O [49], SnapKV [21], LOOK-M [35], and MEDA [34]. We conduct comparisons across different image lengths, different text lengths, and different tasks. KV Cache retention ratio is setting as 0.05. Experimental results indicate that under high compression ratios, FlashCache preserves more of the model’s inherent performance characteristics. At the same time, we find that FlashCache enables the model to answer questions it previously could not. We believe FlashCache filters out noise in some input data, thereby improving the model’s performance under certain conditions.
Appendix G Benchmarks
G.1 Multi-Images Understanding Benchmarks
@@
MileBench[29]. MileBench is a multimodal long-context benchmark designed to evaluate how well Multimodal Large Language Models handle complex multi-image inputs over extended contexts. It is built from 6,440 long-context samples collected from 29 publicly available or self-constructed datasets, with each sample containing multiple images and accompanying text. In our paper, we use four task types from MileBench: Task T (Temporal Multi-Image), which focuses on understanding and reasoning over sequences of images with temporal relationships; Task S (Semantic Multi-Image), which requires integrating complementary semantic information distributed across multiple images; NH (Needle in a Haystack), which tests the model’s ability to retrieve a small but crucial visual or textual detail from long, cluttered multimodal contexts; and IR (Image Retrieval), which assesses whether the model can accurately locate or identify the correct image given a textual description or multimodal query. Together, these tasks make MileBench a comprehensive tool for probing long-context multi-image comprehension, reasoning, and retrieval abilities.
@@
MUIRBench[36]. MUIRBench is a comprehensive benchmark for robust multi-image understanding in multimodal large language models. It contains 11,264 images and 2,600 multiple-choice questions, covering 12 diverse multi-image understanding tasks (such as scene understanding, temporal ordering, and visual retrieval) and 10 categories of inter-image relations (including multiview, narrative, and complementary relations). Each example is constructed in a pairwise manner, where an answerable instance is paired with a minimally modified unanswerable variant, enabling robust evaluation of whether models truly integrate information across multiple images rather than relying on superficial cues. This design makes MUIRBench a strong stress test for multi-image reasoning and a useful complement to standard single-image vision–language benchmarks.
@@
MMMU[46]. MMMU (Massive Multi-discipline Multimodal Understanding) is a large-scale benchmark designed to evaluate foundation models on expert-level, university and beyond problems across a wide range of real-world disciplines. It contains over 10,000 multimodal multiple-choice questions spanning fields such as medicine, law, physics, chemistry, engineering, and the humanities, with inputs that combine text, diagrams, charts, tables, and other images. Many questions involve not just a single image but multiple related figures or panels (for example several diagrams, tables, or subplots for one problem) that must be interpreted jointly, allowing the benchmark to test genuine multi-image reasoning rather than isolated image understanding. All items are carefully curated from real exam or professional materials, emphasizing high-level reasoning, domain knowledge, and cross-modal comprehension, and are widely used to assess how close multimodal large language models are to human expert performance in complex, knowledge-intensive scenarios.
G.2 High Resolution Benchmarks
@@
V*[39]. V* is an LLM-guided visual search mechanism designed to help multimodal large language models actively locate task-relevant regions in an image instead of passively encoding the whole frame. When paired with an MLLM, it underpins the SEAL (“Show, Search, and Tell”) meta-architecture, which iteratively queries the image to refine visual grounding and reasoning. To evaluate this capability, the authors introduce V* Bench, a dedicated benchmark built from 191 high-resolution natural images (average resolution around 2246×1582) that emphasize fine-grained details and visually crowded scenes. V* Bench focuses on detailed visual grounding through attribute recognition and spatial relationship reasoning questions, making it particularly suitable for testing whether models can reliably handle high-resolution inputs and answer questions about small or easily overlooked visual elements.
@@
HR-Bench[38]. HR-Bench is a high-resolution visual question answering benchmark designed to assess the fine-grained perception ability of multimodal large language models on ultra high-resolution images. It consists of two sub-tasks: Fine-grained Single-instance Perception (FSP), which contains 100 samples focusing on detailed attribute recognition, OCR, and visually grounded prompting within a single image, and Fine-grained Cross-instance Perception (FCP), which contains 100 samples targeting map understanding, chart analysis, and spatial relationship reasoning across multiple visual elements. HR-Bench is provided in both 8K and 4K settings: the 8K version uses images with approximately 8K-level resolution, while the 4K version is obtained by carefully annotating objects in the 8K images and cropping them into 4K regions, enabling rigorous evaluation of whether models can maintain and exploit fine details under realistic high-resolution conditions.
G.3 Video Understanding Benchmark
@@
FAVOR-Bench[32]. FAVOR-Bench is a comprehensive benchmark for fine-grained video motion understanding that targets the ability of multimodal large language models to perceive and reason about detailed temporal dynamics rather than just coarse scene changes. It contains 1,776 videos from both ego-centric and third-person perspectives, each accompanied by structured manual motion annotations. On top of this data, FAVOR-Bench provides both close-ended and open-ended evaluation settings: for close-ended evaluation, it offers 8,184 multiple-choice question–answer pairs organized into six motion-centric sub-tasks, while for open-ended evaluation it includes a GPT-assisted caption assessment protocol together with a cost-efficient LLM-free evaluation method. Through these tasks, FAVOR-Bench enables systematic measurement of how well models recognize, localize, and describe fine-grained motions in realistic videos, and serves as a challenging testbed for improving motion-sensitive video understanding.
Appendix H Models
H.1 Qwen2.5-VL-7B-Instruct
@@ Appendix H Models
Qwen2.5-VL-7B-Instruct [37] is a 7-billion-parameter vision–language model from the Qwen2.5-VL family, instruction-tuned to handle a wide range of multimodal tasks with text, images, documents, charts, and video as input. It combines a strong language backbone with a visual encoder that supports dynamic resolutions and long-context processing, enabling capabilities such as document analysis, OCR, chart and layout understanding, temporal reasoning over videos, and structured output generation (for example JSON and bounding boxes for visual localization).
H.2 LLaVA-OneVision-1.5-8B-Instruct
@@
LLaVA-OneVision-1.5-8B-Instruct [4] is an 8-billion-parameter large multimodal model in the LLaVA-OneVision-1.5 family, designed as a fully open, efficient framework for high-quality multimodal training. It is trained on native-resolution images and curated instruction data, aiming to deliver strong performance on diverse vision–language benchmarks while keeping computational costs relatively low. The model follows the LLaVA-style visual instruction tuning paradigm and is positioned as a general-purpose multimodal assistant capable of fine-grained image understanding, reasoning, and dialogue under an open-source training and evaluation pipeline.
Appendix I Limitation and Future Work